# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

### v2

@@ 5.3 Human Evaluation
Method Quality
Merging Captions (Base) 80.5
Merging Videos (Base) 86.5
TALC (Base) 84.5
Merging Captions (F.T.) 63.4
TALC (F.T.) 83.3
@@
We compare the performance of the baselines and TALC framework for ModelScope using human evaluation in Figure 6. We find that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score. In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score. Further, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model. In addition, we find that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively. Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation.
Visual quality of the generated videos.
We compare the visual quality of the generated videos using human evaluation in Table 1. We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework. However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points. This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video. Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining. While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §I.
@@ 5.4 Qualitative Analysis
We provide qualitative examples of generating multi-scene videos using ModelScope (fine-tuned with TALC) and Lumiere (base model with TALC) for diverse scenarios in Figure 12. Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency.
Considering the case of the same animal engaging in multiple actions (referred to as "one character multiple contexts"). The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes. Conversely, the videos generated by Lumiere displayed the same animal performing different actions with minimal background alterations. We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC. Considering different animals within a particular visual setting (referred to as "multiple-characters same context"), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text. During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as "open-ended captions") exhibited a substantial number of changes between the various scenes. In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
@@ 6 Related Work
7 Conclusion
8 Acknowledgement
Appendix A Prompt for Multi-Scene Caption Generation
@@
Appendix B Task Prompts
B.1 Single character in multiple visual contexts
@@
B.2 Different characters in a specific visual context
@@
Appendix C Automatic Evaluation Prompt
@@
Appendix D Human Evaluation Screenshot
@@
Appendix E Fine-grained Analysis
We present the automatic evaluation results across task prompts (§E.1) and number of scenes (§E.2).
E.1 Task Prompts
@@ E.2 Number of Scenes
Appendix F Inference Details
@@
Appendix G Multi-Scene Data Statistics
@@
Appendix H Finetuning Details
@@
Appendix I Single-Scene Video Generation

### v3

@@ 5.2 Human Evaluation Results
We compare the performance of the baselines and TALC framework for ModelScope using human evaluation in Figure 5. We find that TALC-finetuned model outperforms the merging captions and merging video methods with the base model by $12$ points and $15.5$ points, respectively, on the overall score. In addition, we find that using TALC framework in the base model outperforms the merging captions and merging video methods with the base model by $7.6$ points and $11.1$ points, respectively, on the overall score. Further, we observe that the merging captions with the base model achieves the highest visual consistency score of $96.5$ points while it is the lowest for merging videos generated from the base model. In addition, we find that the text adherence of the TALC-finetuned and TALC-base model is better than merging captions-finetuned and merging captions-base model, respectively. Our results highlight at the benefit of including the inductive bias of temporal alignment between the video scenes and their scene descriptions for multi-scene video generation. We perform human evaluation on the visual quality of the generated multi-scene videos in §M. While our work is centered around multi-scene evaluation, we also perform single-scene evaluation in Appendix §N.
@@ 5.3 Qualitative Analysis
We provide qualitative examples of generating multi-scene videos using ModelScope (fine-tuned with TALC) and Lumiere (base model with TALC) for diverse scenarios in Figure 12. Our analysis reveals that both ModelScope and Lumiere are capable of producing multi-scene videos that exhibit high text adherence and visual consistency. Considering the case of the same animal engaging in multiple actions (referred to as ‘one character multiple contexts’). The videos generated by ModelScope successfully maintained the same animal while varying the background and action between the scenes. Conversely, the videos generated by Lumiere displayed the same animal performing different actions with minimal background alterations. We believe that this distinction is attributed to ModelScope’s fine-tuning with TALC. Considering different animals within a particular visual setting (referred to as ‘multiple-characters same context’), both ModelScope and Lumiere demonstrated impressive abilities in preserving the consistency of the background across the videos and adhering closely to the provided text. During our analysis, we noticed that the multi-scene captions derived from real videos (referred to as ‘open-ended captions’) exhibited a substantial number of changes between the various scenes. In this scenario, Lumiere, when employed without fine-tuning, displayed challenges in adhering to the text, while ModelScope achieved a higher degree of text adherence but was also prone to visual artifacts.
@@ 6 Related Work
Text-to-Video Generative Modeling.
The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos. The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis. Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains. Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
Multi-Scene Video Generation.
The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively. Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19]. TALC tackles this task and offers a framework produces videos spanning multiple scenes. We also introduce nuanced evaluation approach. This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis. We provide a detailed related work in §B.
7 Conclusion
8 Acknowledgement
Appendix A Limitations
While there are numerous text-to-video generative models, our work focuses on diffusion-based T2V models due to their higher visual quality and better latency in comparison to autoregressive models [19]. In our work, we argue that the text conditioning mechanism is unified across diffusion T2V models, and thus TALC framework should generalize to various model architectures. Specifically, we perform our experiments with ModelScope, a latent diffusion model, and Lumiere, a diffusion model that operates in the raw video space. In the future, it will helpful to adopt TALC for multi-scene video generation in novel training and model design paradigms.
In this work, we evaluate the multi-scene video generation capabilities uptill four scenes. This is at par with the number of scenes in prior multi-scene video generation works [18, 23]. In our experiments, we further show that the performance of the models degrade as we generate more scenes, and hence, we limited ourselves to four scenes. A relevant future direction would be to expand the quality of multi-scene video-text data beyond four scenes and finetune T2V models using TALC framework.
In addition, we perform human annotations using Amazon Mechanical Turkers (AMT) where most of the workers primarily belong to the US and Canada (hourly rate - \$$\$18$ per hour). Hence, the human annotations in this work do not represent the diverse demographics around the globe. As a result, our human annotations reflect the perceptual biases of the annotators from Western cultures. In the future, it will be pertinent to assess the impact of diverse groups on our human evaluations. Finally, we acknowledge that the text-to-video generative models can perpetrate societal biases in their generated content [45, 46]. It is critical that future work quantifies this bias in the generated videos and provides methods for safe deployment of our models.
Appendix B Detailed Related Work
@@
Appendix C Qualitative comparison between baselines and TALC
We provide an qualitative example to showcase the benefit of TALC over merging captions and merging videos baselines in Figure 6.
Figure 6: Examples of multi-scene video generation baselines. (a) Generating video on the merged descriptions, leads to a poor text-video alignment. (b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of visual consistency. (c) Our approach (TALC) enhances the scene-level text-video alignment and maintains visual consistency.
Appendix D Prompt for Multi-Scene Caption Generation
@@
Appendix E Source datasets for multi-scene video-text data creation
To construct a multi-scene video-text dataset, we utilize existing dataset that include natural (real) videos and associated high-quality human-written captions that summarize the entire video. Specifically, we choose MSR-VTT [53] and VaTeX [54]. Most of the videos in MSR-VTT are 10-30 seconds long while VaTeX consists 10 seconds long videos. In addition, each video in MSR-VTT and VaTex consists 20 captions and 10 captions, respectively, out of which one is selected at random for multi-scene data generation. As described above, a single video is cut into multiple video segments using Pyscene library. In our experiments, we retain the first four video segments and discard any additional segments if the library generates more than four. Since the accuracy of the multi-scene captioning and the computational demands during finetuning are influenced by the number of scenes, we opt to limit the scene count to four for our experiments. However, future work could employ similar methodologies to scale the number of scenes, given more computing power and advanced multi-scene captioning models. We provide the data statistics for the final multi-scene data in Appendix §K.
Appendix F Task Prompts
F.1 Single character in multiple visual contexts
@@
F.2 Different characters in a specific visual context
@@
Appendix G Automatic Evaluation Prompt
@@
Appendix H Human Evaluation Screenshot
@@
Appendix I Fine-grained Analysis
We present the automatic evaluation results across task prompts (§I.1) and number of scenes (§I.2).
I.1 Task Prompts
@@ I.2 Number of Scenes
Appendix J Inference Details
@@
Appendix K Multi-Scene Data Statistics
@@
Appendix L Finetuning Details
@@
Appendix M Visual quality of the generated videos.
Method Quality
Merging Captions (Base) 80.5
Merging Videos (Base) 86.5
TALC (Base) 84.5
Merging Captions (F.T.) 63.4
TALC (F.T.) 83.3
We compare the visual quality of the generated videos using human evaluation in Table 7. We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework. However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points. This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video. Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
Appendix N Single-Scene Video Generation